OpenStudy (anonymous):

Which expressions are equal to 4 when x = 7 and y = 2? Choose all answers that are correct. A. x – y – 1 B. x – (6 ÷ y) C. xy – 8 D. x + y – 4

OpenStudy (anonymous):

@BritBratt13

OpenStudy (mokeira):

substitute

OpenStudy (mokeira):

\[x-y-1= 7-2-1=4\] \[x-(6/y)=7-3=4\] \[xy-8=14-8=6\] \[x+y-4=7+2-4=5\]

OpenStudy (solomonzelman):

plug in 7 for x, and 2 for y, and find out. \(\normalsize\color{blue}{ x-y-1~~~~~⇒ ~~~~~\color{red}{7}-\color{darkgoldenrod}{2} -1}\) \(\normalsize\color{blue}{ x-(6\div y)~~~~~⇒ ~~~~~\color{red}{7}-(6\div \color{darkgoldenrod}{2} )}\) \(\normalsize\color{blue}{ xy-8~~~~~⇒ ~~~~~\color{red}{7}\times \color{darkgoldenrod}{2} -8}\) \(\normalsize\color{blue}{ x+y-4~~~~~⇒ ~~~~~\color{red}{7}+ \color{darkgoldenrod}{2} -4}\)

OpenStudy (solomonzelman):

btw it is called "for all that applies" (or something like this)...

OpenStudy (anonymous):

4 4 6 5 these are the numbers i got

OpenStudy (solomonzelman):

Yes. And you have to have the one(s) that is equal to 4. Which one(s) is it ?

OpenStudy (solomonzelman):

yes, the first 2, A and B .

OpenStudy (mokeira):

@kira1502 thats right. @SolomonZelman has answered :)

OpenStudy (anonymous):

ok thanks so much

OpenStudy (anonymous):

i would medal you both but i can't :/

OpenStudy (solomonzelman):

Just the fact that I have answered doesn't guarantee anything. You are having a false impression because of my purple title :)

OpenStudy (solomonzelman):

medal Mokeira, I have over 5K of medals.

OpenStudy (anonymous):

ok :)

OpenStudy (solomonzelman):

Over 6K :)

OpenStudy (solomonzelman):

have fun !

OpenStudy (anonymous):

ty :)